THE PURE ROTATIONAL SPECTRUM OF FeCO^{+} (X^{4}Σ^{−})

\underline{D. T. HALFEN} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectrum of the molecular ion FeCO^{+} (X^{4}Σ^{−}) has been measured using millimeter-wave direct absorption methods incorporating velocity modulation techniques. This work is the first gas-phase study of FeCO^{+} by any spectroscopic method. FeCO^{+} was created in an AC glow discharge of gas-phase Fe(CO)_{5} and argon. Eleven rotational transitions were measured in the frequency range of 323 to 408 GHz. The four fine structure components exhibited very large spin splittings, indicating a high degree of Hund’s case (a) character and large second-order spin-orbit interactions. The data has been fit with a case (b) Hamiltonian and rotational and fine structure constants were obtained. FeCO^{+} may be prevalent in interstellar/circumstellar gas given the large abundance of CO.